OpenStudy (anonymous):

Please show me the steps to better understand, thank you!!! SOLVE FOR X: 2x+3y=12

OpenStudy (anonymous):

first move y varialbe to other side which will be 2x=12-3y than divide into 2 each side than x=12-3y over 2

OpenStudy (anonymous):

I have an easier way.

OpenStudy (anonymous):

ok send it?

OpenStudy (anonymous):

to find x set y equal to zero. so do 2x + 3 (0) = 12. That way x=6. Then to find y , set x equal to zero. so do 2(0) +3y= 12 so y=4. Done and easy.

OpenStudy (anonymous):

i got 6 and 4 by dividing it from both sides btw. let me know if you have questions

OpenStudy (anonymous):

u need to find x

OpenStudy (anonymous):

x cannot equal to a number

OpenStudy (anonymous):

I'm a bit confused

OpenStudy (anonymous):

please state the steps again please

OpenStudy (anonymous):

umm the point here is to find x right? so u move y to other side which is 3y so it becomes 2x=12-3y and than basically you want to make x alone so u divide 2 by each side than you get x=(12-3y)/2

OpenStudy (anonymous):